# Planning without Search: Refining Frontier LLMs with Offline Goal-Conditioned RL

### v1

@@ 5.2 Evaluated Methods
For each task, we run PNLC with $n\!=\!4$ goals and a single round $m\!=\!2$ of refinement. We chose $n\!=\!4$ to have a mixture containing $2$ positive and negative goals. We evaluate against a range of fine-tuning and prompting baselines; some of these prior methods can be evaluated on the full suite of benchmark tasks, while others are designed specifically for one of the domains. The algorithms can be grouped into the following $3$ classes:
RL fine-tuning. To compare to a prior method that uses multi-turn RL, we use two variants of the recently proposed ArCHeR algorithm [48], which uses an actor-critic architecture to directly fine-tune LLMs for interactive tasks. We compare to two variants of a popular fine-tuning algorithm: (1) offline ArCHer and (2) online ArCHer, which employs interactive rollouts through the environment to further optimize the agent’s behavior. Due to the complexity of end-to-end value-based RL methods, we run ArCHeR using the smaller GPT-2 model [28] to represent the policy, following prior work [48].
LLM reasoning. Next, we evaluate a variety of sophisticated strategies used to invoke reasoning and planning capabilities on a GPT-4 LLM agent [22], the most complex of which performs inference-time search. We compare to: (1) ReAct [41], which uses chain-of-thought prompting for LLM agents to reason about their actions solely using the zero-shot capabilities of the LLM; (2) Reflexion  [31], which adds $n\!=\!5$ additional simulations of full trajectories that the LLM agent is allowed to refine with; (3) LATS  [47], which performs full MCTS simulation and selects its actions based on the results of the search. We evaluate both a fast and slow version of each method, with either $n\!=\!5$ or $n\!=\!30$ simulations, to provide a fairer point of comparison for our method (the fast version requires a similar inference budget as ours). Finally, we evaluate a custom adaptation of Huang et al. [9] from question-answering to multi-turn tasks called (4) Self-Plan; rather than using majority vote on $n$ answers by the model, we instead have the model select from $n\!=\!5$ of its own generations. In contrast to LATS, Self-Plan does not perform any search, but performs self-refinement by assessing its own generations.
@@
    WebShop: Putta et al. [26] achieved state-of-the-art performance with their method Agent Q, which mixes offline fine-tuning with inference-time search. In the offline phase, rather than value-based RL, Agent Q uses DPO  [29] on preference pairs of trajectories in the dataset. Then, during inference, Agent Q performs MCTS search to generate $n$ simulations before selecting an environment action. Like with LATS, we consider fast $n\!=\!5$ and slow $n\!=\!30$ variations of Agent Q. Note that because Agent Q requires modifying the weights of the base LLM agent, they train on Mixtral-7B [10] as more powerful models such as GPT-4 [22] only expose inference APIs.
@@
    AvalonBench: Strategist [17] proposes a hierarchical prompting mechanism that additionally uses MCTS via population-based self-play to propose and refine high-level strategies. The high-level strategies searched over by Strategist are similar to the thoughts refined by our proposed method, but ours does not rely on any self-play during inference. Again, to keep a fair comparison with inference time, we consider both a fast $n\!=\!5$ and slow $n\!=\!30$ variations of Strategist varying the number of simulations during search.
@@
    Persuasion: Yu et al. [44] propose GDP-Zero, which prompts the LLM to perform tree-search over possible high-level strategies at every timestep, simulating responses by both interlocutors in the dialogue, then selects the best action according to the search. This method is similar to LATS, but searches over a task-specific persuasion taxonomy rather than the full space of actions. We consider $n\!=\!5$ and $n\!=\!30$ varations of the approach, varying the number of simulated dialogues.
@@
ArCHer  [48] $62.3$ $32.0$ $19.0$ $0.36$
Offline ArCHer  [48] $57.3$ $28.0$ $18.0$ $0.31$
ReAct  [41] $55.1$ $27.0$ $21.0$ $0.54$
Reflexion  [31] $60.8$ $29.0$ $26.0$ $0.54$
LATS ($n\!=\!30$)  [47] $74.9$ $44.0$ $38.0$ $0.78$
LATS ($n\!=\!5$)  [47] $53.9$ $28.0$ $22.0$ $0.52$
Self-Plan  [9] $54.8$ $28.0$ $27.0$ $0.55$
Agent Q ($n\!=\!30$)  [26] $77.1$ $\mathbf{48.0}$
Agent Q ($n\!=\!5$)  [26] $63.2$ $35.0$
Strategist ($n\!=\!30$)  [17] - - $42.0$
Strategist ($n\!=\!5$[17] - - $31.0$
GDP-Zero ($n\!=\!30$)  [44] - - - $0.74$
GDP-Zero ($n\!=\!5$)  [44] - - - $0.47$
@@ 5.3 Results
A table of all results can be found for all considered benchmarks in Table 1. We see that PNLC performs best across all tasks, the closest competitors being task-specific approaches such as Agent Q in WebShop or Strategist in Avalon. However, it is important to note that such methods use domain-specific prompting mechanisms and do not naively generalize to other tasks; in addition, Agent Q requires DPO fine-tuning on a Mixtral-7B LLM, which is much more expensive than the lightweight value function that we train. Furthermore, the $n\!=\!30$ version of search approaches requires around 10×10$\times$10 × more time to make a single decision — $46$s for Agent Q compared to $5$s for ours in WebShop, and $62$s for Strategist compared to $6$s for ours in Avalon. Interestingly, though RL fine-tuning approaches such as ArCHer train on the same data as ours, but actually often perform the worst. We attribute this to the fact that standard RL fine-tuning methods rely on a much weaker LLM as the agent; qualitatively, such LLM agent often gives nonsensical or off-topic responses.
@@ 5.4 Ablation Study
Finally, we conducted an ablation analysis to better understand where the improvements in our method were coming from. We test two important components of our method: (1) offline goal-conditioned training to obtain natural language values, and (2) a novel inference procedure of refining thoughts using these values.
To evaluate the efficacy of each component, we evaluate two modifications of PNLC. The first removes the goal-conditioning, instead learning a scalar Q-value conditioned only state and thought, which simply predicts likelihood of success in the task. Recall that we use goals in our natural language critic in order to create informative descriptions that are more naturally used during LLM reasoning. Now, the LLM agent sees a scalar value instead of natural language, and makes its refinement off that value. The second modification replaces the inference-time refinement procedure with more traditional best-of-$n$ selection [15, 30]. Instead of being given the chance to refine its original thought using feedback provided by the natural language critic, this method instead has the LLM agent generate $n\!=\!5$ thoughts, from which one is chosen by the same agent. Note that our refinement procedure was designed with the hypothesis that an LLM agent can produce better thoughts after observing what may go wrong, which this modification cannot leverage.
We show results in Table 2. We notice that the two ablations perform comparably to other LLM reasoning baselines such as ReAct that rely only on prompting, meaning the LLM agent does not effectively use the auxiliary learned value function. This shows that both components of our approach — the ability to (1) craft natural language descriptions via goals and (3) refine thoughts using such descriptions — are both crucial to the success of our algorithm.
@@
ReAct  [41] $55.1$ $27.0$ $21.0$ $0.54$
PNLC w/o goals $55.4$ $27.0$ $25.0$ $0.53$
PNLC w/o refinement $55.6$ $30.0$ $28.0$ $0.61$
@@
Table 2: Mean evaluation results for ablations across 100 independent test runs. We see that both ablations of our method perform much worse, meaning that both the goal-conditioning and self-refinement process are important to our method. Without both, the ablations match the performance of simple chain-of-thought prompting.
@@ 6 Discussion
Limitations. The most notable limitation of our approach is the need to train task-specific value functions. Though our value functions are light-weight and easy to train, an important direction of future research is the adaptation of our method to allow generalist reasoning in LLM agents. Furthermore, another limitation of our method is the reliance on an LLM’s intuition to reason about possible future states. This may be hard to do for tasks that are out-of-domain of an LLM’s base reasoning capabilities. For example, it may be risky to rely on an LLM’s “intuition” for tasks like medical diagnosis, which require a plethora of domain-specific knowledge. Finally, another important avenue of future research is an investigation of the quality and diversity of the offline data necessary for our approach to work well. In our work, we naïvely prompt an LLM to behave as the data collection policy; it is unclear if such policy will prove effective across all tasks, particularly ones where naive LLM agents perform very poorly.
Broader impacts. Our work attempts to make contributions in the development of capable LLM agents for interactive tasks, including dialogue. Superhuman abilities in the realm of persuasion can be used for harmful applications, and our work does not at this time provide any mechanism for avoiding or mitigating such harms. Our method is one of many algorithms recently developed to learn LLM agents, which are inherently at risk for dual use. The focus of our paper is on improving scalability, and we evaluate only on publicly available testbeds.
@@ A.3 Avalon Prompts
For game rules, we use the prompt specified in the Appendix of Light et al. [16]. Similarly, for generating actions for proposing teams (when chosen as leader), voting, or for discussion, we use the prompts used by Strategist as specified in the Appendix of Light et al. [17].

### v2

@@ 5.2 Evaluated Methods
For each task, we run PNLC with n=4n\!=\!4 goals and a single round m=2m\!=\!2 of refinement. We chose n=4n\!=\!4 to have a mixture containing $2$ positive and negative goals. We evaluate against a range of fine-tuning and prompting baselines; some of these prior methods can be evaluated on the full suite of benchmark tasks, while others are designed specifically for one of the domains. The algorithms can be grouped into the following $3$ classes:
RL fine-tuning. To compare to a prior method that uses multi-turn RL, we use two variants of the recently proposed ArCHeR algorithm [zhou2024archertraininglanguagemodel], which uses an actor-critic architecture to directly fine-tune LLMs for interactive tasks. We compare to two variants of a popular fine-tuning algorithm: (1) offline ArCHer and (2) online ArCHer, which employs interactive rollouts through the environment to further optimize the agent’s behavior. Due to the complexity of end-to-end value-based RL methods, we run ArCHeR using the smaller GPT-2 model [radford2019language] to represent the policy, following prior work [zhou2024archertraininglanguagemodel].
LLM reasoning. Next, we evaluate a variety of sophisticated strategies used to invoke reasoning and planning capabilities on a GPT-4 LLM agent [OpenAI2023GPT4], the most complex of which performs inference-time search. We compare to: (1) ReAct [yao2022react], which uses chain-of-thought prompting for LLM agents to reason about their actions solely using the zero-shot capabilities of the LLM; (2) Reflexion  [shinn2023reflexionlanguageagentsverbal], which adds n=5n\!=\!5 additional simulations of full trajectories that the LLM agent is allowed to refine with; (3) LATS  [zhou2024languageagenttreesearch], which performs full MCTS simulation and selects its actions based on the results of the search. We evaluate both a fast and slow version of each method, with either n=5n\!=\!5 or n=30n\!=\!30 simulations, to provide a fairer point of comparison for our method (the fast version requires a similar inference budget as ours). Finally, we evaluate a custom adaptation of huang2023improve from question-answering to multi-turn tasks called (4) Self-Plan; rather than using majority vote on $n$ answers by the model, we instead have the model select from n=5n\!=\!5 of its own generations. In contrast to LATS, Self-Plan does not perform any search, but performs self-refinement by assessing its own generations.
@@
    WebShop: putta2024agentqadvancedreasoning achieved state-of-the-art performance with their method Agent Q, which mixes offline fine-tuning with inference-time search. In the offline phase, rather than value-based RL, Agent Q uses DPO  [rafailov2023direct] on preference pairs of trajectories in the dataset. Then, during inference, Agent Q performs MCTS search to generate $n$ simulations before selecting an environment action. Like with LATS, we consider fast n=5n\!=\!5 and slow n=30n\!=\!30 variations of Agent Q. Note that because Agent Q requires modifying the weights of the base LLM agent, they train on Mixtral-7B [jiang2024mixtralexperts] as more powerful models such as GPT-4 [OpenAI2023GPT4] only expose inference APIs.
@@
    AvalonBench: Strategist [light2024strategistlearningstrategicskills] proposes a hierarchical prompting mechanism that additionally uses MCTS via population-based self-play to propose and refine high-level strategies. The high-level strategies searched over by Strategist are similar to the thoughts refined by our proposed method, but ours does not rely on any self-play during inference. Again, to keep a fair comparison with inference time, we consider both a fast n=5n\!=\!5 and slow n=30n\!=\!30 variations of Strategist varying the number of simulations during search.
@@
    Persuasion: yu2023promptbasedmontecarlotreesearch propose GDP-Zero, which prompts the LLM to perform tree-search over possible high-level strategies at every timestep, simulating responses by both interlocutors in the dialogue, then selects the best action according to the search. This method is similar to LATS, but searches over a task-specific persuasion taxonomy rather than the full space of actions. We consider n=5n\!=\!5 and n=30n\!=\!30 varations of the approach, varying the number of simulated dialogues.
@@
ArCHer  [zhou2024archertraininglanguagemodel] $62.3$ $32.0$ $19.0$ $0.36$
Offline ArCHer  [zhou2024archertraininglanguagemodel] $57.3$ $28.0$ $18.0$ $0.31$
ReAct  [yao2022react] $55.1$ $27.0$ $21.0$ $0.54$
Reflexion  [shinn2023reflexionlanguageagentsverbal] $60.8$ $29.0$ $26.0$ $0.54$
LATS (n=30n\!=\!30)  [zhou2024languageagenttreesearch] $74.9$ $44.0$ $38.0$ $0.78$
LATS (n=5n\!=\!5)  [zhou2024languageagenttreesearch] $53.9$ $28.0$ $22.0$ $0.52$
Self-Plan  [huang2023improve] $54.8$ $28.0$ $27.0$ $0.55$
Agent Q (n=30n\!=\!30)  [putta2024agentqadvancedreasoning] $77.1$ $\mathbf{48.0}$ - -
Agent Q (n=5n\!=\!5)  [putta2024agentqadvancedreasoning] $63.2$ $35.0$ - -
Strategist (n=30n\!=\!30)  [light2024strategistlearningstrategicskills] - - $42.0$ -
Strategist (n=5n\!=\!5[light2024strategistlearningstrategicskills] - - $31.0$ -
GDP-Zero (n=30n\!=\!30)  [yu2023promptbasedmontecarlotreesearch] - - - $0.74$
GDP-Zero (n=5n\!=\!5)  [yu2023promptbasedmontecarlotreesearch] - - - $0.47$
@@ 5.3 Results
A table of all results can be found for all considered benchmarks in Table 1. We see that PNLC performs best across all tasks, the closest competitors being task-specific approaches such as Agent Q in WebShop or Strategist in Avalon. However, it is important to note that such methods use domain-specific prompting mechanisms and do not naively generalize to other tasks; in addition, Agent Q requires DPO fine-tuning on a Mixtral-7B LLM, which is much more expensive than the lightweight value function that we train. Furthermore, the n=30n\!=\!30 version of search approaches requires around $10\times$ more time to make a single decision — $46$s for Agent Q compared to $5$s for ours in WebShop, and $62$s for Strategist compared to $6$s for ours in Avalon. Interestingly, though RL fine-tuning approaches such as ArCHer train on the same data as ours, but actually often perform the worst. We attribute this to the fact that standard RL fine-tuning methods rely on a much weaker LLM as the agent; qualitatively, such LLM agent often gives nonsensical or off-topic responses.
@@ 5.4 Ablation Study
We conducted an ablation analysis to better understand where the improvements in our method were coming from. We test two important components of our method: (1) offline goal-conditioned training to obtain natural language values, and (2) a novel inference procedure of refining thoughts using these values.
To evaluate the efficacy of each component, we evaluate two modifications of PNLC. The first removes the goal-conditioning, instead learning a scalar Q-value conditioned only state and thought, which simply predicts likelihood of success in the task. Recall that we use goals in our natural language critic in order to create informative descriptions that are more naturally used during LLM reasoning. Now, the LLM agent sees a scalar value instead of natural language, and makes its refinement off that value. The second modification replaces the inference-time refinement procedure with more traditional best-of-$n$ selection [li2024QProbe, sessa2024bondaligningllmsbestofn]. Instead of being given the chance to refine its original thought using feedback provided by the natural language critic, this method instead has the LLM agent generate n=5n\!=\!5 thoughts, from which one is chosen by the same agent. Note that our refinement procedure was designed with the hypothesis that an LLM agent can produce better thoughts after observing what may go wrong, which this modification cannot leverage.
Finally, we include an ablation dubbed ReAct+Replan that does not affect the inference pipeline at all, so the same self-refinement inference method using subgoals is preserved, but the scores of subgoals are directly inferred by the frontier LLM rather than learned via offline goal-conditioned RL. This ablation tests whether any training is necessary, or whether the scaffolding we propose can be handled entirely using the base capabilities of frontier LLMs.
We show results in Table 2. We notice that the first two ablations perform comparably to other LLM reasoning baselines such as ReAct that rely only on prompting, meaning the LLM agent does not effectively use the auxiliary learned value function. This shows that both components of our approach — the ability to (1) craft natural language descriptions via goals and (3) refine thoughts using such descriptions — are both crucial to the success of our algorithm. Finally, ReAct+Replan performs better but still noticeably worse than our method, which shows that offline training is important, and that even frontier LLMs are incapable of providing accurate scores for goals in long-horizon tasks a priori. Qualitatively, we see that even frontier LLMs provide uncalibrated, often overly-optimistic scores, such that they rarely deem their own actions to be suboptimal. In this manner, the refinement stage in ReAct+Replan often only decides that the initial action was good.
@@
ReAct  [yao2022react] $55.1$ $27.0$ $21.0$ $0.54$
PNLC (w/o goals) $55.4$ $27.0$ $25.0$ $0.53$
PNLC (w/o refinement) $55.6$ $30.0$ $28.0$ $0.61$
ReAct+Replan $59.1$ $32.0$ $22.0$ $0.62$
@@
Table 2: Mean evaluation results for ablations across 100 independent test runs. We see that all ablations of our method perform much worse, meaning that both the goal-conditioning and self-refinement process are important to our method. Without either, the ablations match the performance of simple chain-of-thought prompting.
@@ 6 Discussion
Limitations. The most notable limitation of our approach is the need to train task-specific value functions. Though our value functions are light-weight and easy to train, an important direction of future research is the adaptation of our method to allow generalist reasoning in LLM agents. Furthermore, another limitation of our method is the reliance on an LLM’s intuition to reason about possible future states. This may be hard to do for tasks that are out-of-domain of an LLM’s base reasoning capabilities. For example, it may be risky to rely on an LLM’s “intuition” for tasks like medical diagnosis, which require a plethora of domain-specific knowledge. Finally, another important avenue of future research is an investigation of the quality and diversity of the offline data necessary for our approach to work well.
@@ A.3 Avalon Prompts
For game rules, we use the prompt specified in the Appendix of light2023avalonbenchevaluatingllmsplaying. Similarly, for generating actions for proposing teams (when chosen as leader), voting, or for discussion, we use the prompts used by Strategist as specified in the Appendix of light2024strategistlearningstrategicskills.